1. The number of Bacteria in a particular colony is found to double every 2.25 seconds. If the initial count of bacteria is 12500, what will be the population after 125 seconds?


2. The distance covered by a freefalling object under the action of constant action of gravitational field quadratically varies with time. In a time of 3.0 seconds, a free-falling object covers a 5.0-meter distance. what distance will it cover in 5.0 seconds?


1
Expert's answer
2022-03-30T17:43:01-0400

1.

"n=n_02^{\frac{t_2}{t_1}}=6.6\cdot 10^{20}."

2.

"s_1=kt_1^2,"

"k=\frac{s_1}{t_1^2}=0.55,"

"s_2=kt^2_2=13.9~m."